# Gauge Symmetry as the Origin of Spin–Orbit Coupling of Light

### v4

@@ I Introduction
$\psi^{\pm}=\mathcal{D}_{\perp}\pm i\sigma_{3}\mathcal{B}_{\perp},$ (3)
@@ II Dual gauge field
According to the principles of gauge theory [45], maintaining local symmetry requires the introduction of a corresponding gauge field. In general, the infinitesimal variation of the derivative term under a local transformation can be expressed as
$\delta(\partial_{i}\psi_{\sigma})=-ig\partial_{i}(\theta(x)\psi_{\sigma}),$ (6)
where $g$ is some coupling coefficient. To preserve gauge invariance and compensate for this additional term, a gauge field $A_{i}$ is introduced, thereby promoting the ordinary derivative to a covariant derivative,
$\partial_{i}\rightarrow D_{i},\quad D_{i}=\partial_{i}+igA_{i},\;\rightarrow\;\hat{K}_{i}=\hat{k}_{i}+gA_{i},$ (7)
where $\hat{K}_{i}$ represents the covariant derivative in the momentum operator form. In this context, the gauge potential $A_{i}$ is associated with the spatial gradient of the material parameter, typically taking the form $\partial_{i}\ln\epsilon^{-1}$.
@@
$\mathcal{L}=\mathcal{L}_{0}+\mathcal{L}_{A},$ (8)
where $\mathcal{L}_{0}$ represents the free photon part associated with global dual symmetry and can be constructed from the corresponding optical Dirac equation [20], while $\mathcal{L}_{A}$ accounts for the interaction between the photon and the gauge field, arising from the localization of the symmetry via $\theta(x)$.
@@
[…]
@@
To preserve gauge invariance, the gauge field must transform in a way that compensates for this coordinate-dependent variation, namely $A_{i}\rightarrow A_{i}+\partial_{i}\theta$. By comparing this result with the covariant derivative defined in Eq. (7), the coupling coefficient $g$ can be identified with the matrix $\gamma^{0}\hat{S}_{z}$, which corresponds to the helicity operator in the helicity eigenfunction representation (see Sec.III).
If we retain only the leading term, the interaction between the gauge field and the photon field can be expressed as
$\mathcal{L}_{A}=\overline{\psi}\{-\gamma^{0}\hat{S}_{z}(Q_{0}A_{z}+\bm{q}\cdot\bm{A})+\gamma^{\pm}\hat{S_{z}}(Q_{\pm 2}A_{z}-2Q_{\pm 1}A_{\pm})\}\psi.$ (17)
This result implies that when the medium is decomposed into tensor modes according to the concept of the original excitation, interactions naturally arise between these modes, photons, and the dual gauge fields.
@@
where $\mathcal{L}_{Q}$ captures terms involving the first-order spatial derivatives[46],
@@
When the curl of the dipole moment vanishes—i.e., when $\hat{L}_{q}^{z}=0$—a trivial local gauge field arises. In this case, the gauge potential takes the pure-gauge form $A_{i}=iU^{\dagger}\nabla_{i}U$, where the transformation factor take $U=e^{-i\gamma^{0}\hat{S}_{z}\theta(\mathbf{x})}$. This corresponds to a vanishing gauge field strength, $F_{ij}=0$. By identifying Eqs. (17) and (19), we find that the following condition must hold:
$\hat{k}_{i}Q_{j}=\gamma^{0}\hat{S}_{z}Q_{j}A_{i},$ (23)
@@
$\triangle Q_{j}\approx\lambda A_{i}\triangle x^{i},$ (25)
@@
In the following chapter, we will examine how the spatial rotation of the dipole term $\bm{q}$ along the helical trajectories of photons induces nontrivial gauge fields, and show that these dipole-driven effects constitute the physical origin of the gauge fields associated with spin–orbit coupling.
III Gauge field representation of spin-orbit coupling
Since the helical eigenstate of light serves as an independent propagation quantity under spin-degenerate conditions, we can formulate the $U(1)$ dual symmetry by taking the helical eigenstate as the fundamental element. Let $\psi_{i}$ denote the $i$th-component of a four-vector field $\psi$. The constraint imposed by helicity conservation allows the wavefunction to be decomposed into two opposite helicity states $\psi_{\sigma}$ with the helicities $\lambda=\pm 1$: $\psi_{+1}=(\psi_{1},0,0,\psi_{4})^{T}$, and $\psi_{-1}=(0,\psi_{2},\psi_{3},0)^{T}$. In this helicity representation, the coupling term $\gamma^{0}\hat{S}_{z}$ can be interpreted directly in terms of the helicity eigenvalue $\lambda$. Under a local dual transformation, the helicity component $\psi_{\lambda}$ transforms as
@@ III Gauge field representation of spin-orbit coupling
The covariant derivative $D_{0}$ with respect to time can equivalently be expressed in terms of the path length derivative $\hat{D}_{s}$. In this representation, the only nonvanishing component of the gauge field appears along the propagation path, yielding
$D_{s}=\partial_{s}+i\gamma^{0}\hat{S}_{z}{\mathcal{A}}_{s}.$ (28)
Obviously, under the local transformation introduced in Eq. (4), incorporating the gauge field transforms,
@@
the Lagrangian keeps invariant. Accordingly, the effective Lagrangian associated with the gauge field component $A_{s}$ in the helicity eigenfunction representation $\psi_{\lambda}$ is
@@
Since photons can also carry intrinsic orbital angular momentum (IOAM), $\hat{\bf L}=\hat{\bf x}\times\hat{\bf k}$, we generalize the photon wave function to a twisted optical field, written as $\psi^{+}e^{il\varphi}$ and $\psi^{-}e^{-il\varphi}$, where $l$ is the topological charge and $\varphi$ is the azimuthal angle [2]. For such a twisted light beam, an additional first-order correction term arises from the coupling between IOAM and EOAM, $\sim\hat{\bf L}\cdot\hat{\bf L}_{q}$, which contributes to the Lagrangian $\mathcal{L}_{0}$ at the same order as Eq. (31):
@@
which, for a vortex beam with topological charge $l$, becomes $\lambda_{\sigma}=(l+\sigma)$, where $\sigma=\pm 1$ denotes the spin eigenvalue of the photon.
A corresponding IOAM–EOAM coupling term also arises in the Lagrangian $\mathcal{L}_{A}$. To relate this term to the gauge field, however, the original dual symmetry must be generalized:
$\psi_{\lambda_{\sigma}}\rightarrow e^{-i\lambda_{\sigma}\theta(s)}\psi_{\lambda_{\sigma}}.$ (36)
This transformation represents a local rotation that acts not only on the polarization basis but also on the wavefront structure, effectively shifting the azimuthal phase from $\varphi$ to $\varphi-\theta(s)$. For the Lagrangian in Eq. (27) to remain invariant under this local symmetry transformation, the medium response $\delta Q_{i}$ must compensate for the non-vanishing commutator between the OAM operator $\hat{L}_{z}$ and the Lagrangian. Physically, this variation describes the angular momentum transfer between the photon and the medium, potentially incorporating contributions from induced multipole moments. Consequently, reassigning a term from $\mathcal{L}_{0}$ to $\mathcal{L}_{A}$ requires the introduction of a compensating contribution that accounts for the medium’s variation.
The interaction between the gauge field and the photon wavefunction in momentum representation can be expressed as
$\mathcal{L}_{A}=-\sum_{\sigma}\lambda_{\sigma}\psi_{\lambda_{\sigma}}^{\dagger}\dot{\mathbf{k}}\cdot\mathbf{A}\psi_{\lambda_{\sigma}}.$ (37)
which is consistent with the definition of angular velocity in momentum space, $\Omega(s)=-\mathbf{A}\cdot\dot{\mathbf{k}}$ [8]. This formalism naturally extends to momentum space through the introduction of a covariant derivative,
$aligned D_{s}^{\sigma}=\partial_{s}+i\lambda_{\sigma}{\mathcal{A}}_{s}\rightarrow\frac{dk^{i}}{ds}D_{i}^{\sigma}=\frac{dk^{i}}{ds}(\partial_{i}+i(\lambda_{\sigma}{\mathcal{A}}_{i})),$ (38)
Accordingly, the covariant derivative in momentum space is defined as
$D_{i}^{\sigma}=\partial_{i}+i\lambda_{\sigma}{\mathcal{A}}_{i},$ (39)
and its commutator yields
$D_{i}^{\sigma},D_{j}^{\sigma}]=i\lambda_{\sigma}{\mathcal{F}}_{ij}.$ (40)
[…]
From the perspective of the classical analogy to quantum mechanics, Eq. (37) can be interpreted as the spin–orbit interaction term. It can also be viewed as a Coriolis-like effect arising in a non-inertial reference frame co-moving with the wave [9, 8]. The corresponding Euler–Lagrange equations for the classical variables $\mathbf{k}$ and $\mathbf{r}$ reproduce the conventional ray trajectory. The inclusion of spin–orbit coupling introduces a first-order correction to the group velocity, resulting in the optical Hall effect [46, 12]:
[…]
$U_{g}=\exp[i\lambda_{\sigma}\oint_{\Gamma}dk^{i}A_{i}(k)]=\exp[i\lambda_{\sigma}\oint_{\Sigma}dS^{ij}F_{ij}(k)],$ (42)
@@ IV Conclusion
We investigate the localization of the dual symmetry of the electromagnetic field in helical systems using a field-theoretical framework, in which a novel gauge field emerges naturally as a consequence of the local symmetry. The spin–orbit interaction arising during beam propagation along helical trajectories is formally equivalent to the coupling between the helicity and this emergent gauge field. This correspondence provides a unified gauge-theoretic interpretation of the spin Hall effect of light, where such phenomena can be viewed as analogues of axion–electromagnetic coupling [43, 23, 36, 42]. For vortex beams, corrections from orbit–orbit interactions remain significant, yet these are fundamentally distinct from spin–orbit coupling, which originates from a deeper intrinsic symmetry. This distinction provides a fresh perspective for differentiating between the spin and IOAM of light.
A particularly important result is Eq. (30). In conventional electrodynamics, it is well known that charge conservation arises naturally from the U(1)U(1)gauge symmetry of the electronic system, and that promoting this symmetry to a local one introduces the electromagnetic gauge potential—enabling interactions between electrons via photon exchange. In a completely analogous manner, the dual symmetry considered here is also a $U(1)$ symmetry, whose associated conserved quantity is helicity. Localizing this dual symmetry introduces a dual gauge field, implying that photons interact with each other through the exchange of quanta associated with this field. Consequently, our framework may provide a fundamentally new theoretical basis for describing interactions between light and matter.

### v5

@@ I Introduction
$\psi^{\pm}=\mathcal{D}_{\perp}\pm i\sigma_{3}\mathcal{B}_{\perp}.$ (3)
@@ II Dual gauge field
According to the principles of gauge theory  [47], preserving a local symmetry requires the introduction of an associated gauge field. From an optical perspective, the transformation in Eq. (4) may be viewed as a position-dependent rotation in the internal (spin/helicity) space of the field. When the transformation parameter $\theta(\mathbf{x})$ varies in space, taking a spatial derivative necessarily acts on both the field and the parameter, thereby producing an extra term that cannot be removed by the ordinary derivative alone. For an infinitesimal transformation, this reads
$\delta(\partial_{i}\psi)=-i\gamma^{0}\hat{S}_{z}\partial_{i}(\theta(\mathbf{x})\psi).$ (6)
Equivalently, the derivative generates an additional contribution proportional to $\partial_{i}\theta(\mathbf{x})$, i.e., an effective “connection” acting in the helicity (spin) subspace. To preserve gauge invariance and compensate for this extra term, we introduce a dual gauge field $A_{i}$ and promote the ordinary derivative to a covariant derivative,
$\partial_{i}\rightarrow D_{i},\quad D_{i}=\partial_{i}+i\gamma^{0}\hat{S}_{z}\mathcal{A}_{i},\;\rightarrow\;\hat{K}_{i}=\hat{k}_{i}+\gamma^{0}\hat{S}_{z}\mathcal{A}_{i},$ (7)
where $\hat{K}_{i}$ denotes the momentum-operator representation of the covariant derivative. The gauge field $A_{i}$ is determined by, and encodes, variations of the medium. In Sec. III, we derive an explicit representation of the coupling factor $\gamma^{0}\hat{S}_{z}$in terms of the photon helical eigenfunctions.in terms of the photon helical eigenfunctions. In the present context, this coupling constant is associated with the photon helicity and thus corresponds to a conserved quantity protected by dual symmetry, in contrast to the electric charge $e$ that sets the coupling strength in electrodynamics.
@@
$\mathcal{L}=\mathcal{L}_{0}+\mathcal{L}_{\mathcal{A}},$ (8)
where $\mathcal{L}_{0}$ represents the free photon part associated with global dual symmetry and can be constructed from the corresponding optical Dirac equation  [20], while $\mathcal{L}_{\mathcal{A}}$ accounts for the interaction between the photon and the gauge field, arising from the localization of the symmetry via $\theta(x)$.
@@ II.1 Pure Gauge Potential
It can be verified that the system remains invariant under the gauge transformations given in Eq.(4). When a local transformation parameter $\theta(\bm{x})$ is introduced, substituting this transformation into the Lagrangian in Eq.(13) shifts all momentum operators according to
@@
Accordingly, the gauge field transform as $\mathcal{A}_{i}\rightarrow\mathcal{A}_{i}+\partial_{i}\theta$ so as to compensate for the coordinate-dependent variation.
Retaining only the leading-order contribution, the interaction between the gauge field and the photon field can be written as
$\mathcal{L}_{A}=\overline{\psi}\{-\gamma^{0}\hat{S}_{z}(Q_{0}\mathcal{A}_{z}+\bm{q}\cdot\mathbf{\mathcal{A}})+\gamma^{\pm}\hat{S_{z}}(Q_{\pm 2}\mathcal{A}_{z}-2Q_{\pm 1}\mathcal{A}_{\pm})\}\psi.$ (17)
Here, the effects of spatial variations in the material tensor are encoded in the dual gauge field $\mathcal{A}_{i}$. This expression describes an additional light–matter interaction mediated by the elementary excitations in terms of multipole modes, the induced gauge field, and the light field.
@@
where $\mathcal{L}_{Q}$ captures terms involving first-order spatial derivatives [48],
@@
When both the curl and the divergence of the dipole moment in Eq. (20) vanish, the corresponding local gauge field becomes trivial. In this case, the gauge potential takes the pure-gauge form $\mathcal{A}_{i}=iU^{\dagger}\nabla_{i}U$, where the transformation factor take $U=e^{-i\gamma^{0}\hat{S}_{z}\theta(\mathbf{x})}$. This corresponds to a vanishing gauge field strength, $\mathcal{F}_{ij}=0$. By identifying Eqs. (17) and (19), we find that the following condition must hold:
$\hat{k}_{i}Q_{j}=\gamma^{0}\hat{S}_{z}Q_{j}\mathcal{A}_{i},$ (23)
@@
$\triangle Q_{j}\approx\lambda\mathcal{A}_{i}\triangle x^{i},$ (25)
@@
II.2 Nontrivial Dual Gauge Field
Since the helical eigenstate of light serves as an independent propagation quantity under spin-degenerate conditions, we can formulate the $U(1)$ dual symmetry by taking the helical eigenstate as the fundamental element. Let $\psi_{i}$ denote the $i$th-component of a four-vector field $\psi$. The constraint imposed by helicity conservation allows the wavefunction to be decomposed into two opposite helicity states $\psi_{\sigma}$ with the helicities $\lambda=\pm 1$: $\psi_{+1}=(\psi_{1},0,0,\psi_{4})^{T}$, and $\psi_{-1}=(0,\psi_{2},\psi_{3},0)^{T}$. Under a local dual transformation, the helicity component $\psi_{\lambda}$ transforms as
@@ II.2 Nontrivial Dual Gauge Field
In this helicity representation, the coupling term $\gamma^{0}\hat{S}_{z}$ can be expressed directly in terms of the helicity eigenvalue $\lambda$. Furthermore, one can verify that, when Eq. (13) is transformed to this basis, the total Lagrangian becomes block-diagonal, decomposing into two independent $2\times 2$ diagonal blocks. This structure implies that the helicity eigenstates decouple and their corresponding wavefunctions evolve independently.
@@
The covariant derivative $D_{0}$ with respect to time can equivalently be expressed in terms of the path length derivative $D_{s}$.
In the helical representation $\psi_{\lambda}$, the only nonvanishing component of the gauge field appears along the propagation path, yielding
$D_{s}=\partial_{s}+i\lambda{\mathcal{A}}_{s}.$ (28)
Under the local transformation introduced in Eq. (4), the gauge field transforms as
@@
leaving the Lagrangian invariant. The covariant-derivative structure can be naturally extended to momentum space by introducing
$aligned D_{s}=\partial_{s}+i\lambda{\mathcal{A}}_{s}\rightarrow\frac{dk^{i}}{ds}D^{k}_{i}=\frac{dk^{i}}{ds}(\frac{\partial}{\partial k_{i}}+i\lambda{\mathcal{A}}_{i}(k)),$ (30)
or equivalently,
$D_{i}^{k}=\partial^{k}_{i}+i\lambda{\mathcal{A}}_{i}(k).$ (31)
We introduce the following momentum-space representation by making the substitution:
$R_{i}\rightarrow\partial^{k}_{i},\quad r_{i}\rightarrow D^{k}_{i},$ (32)
where $R$ denotes the canonical coordinate and $r$ the physical coordinate. They are related through the gauge field $\bm{A}(k)$ via $\bm{r}=\bm{R}-\bm{A}(k)$ [10, 48]. The Lagrangian associated with the corresponding covariant derivative in Eq. (27) is
$\mathcal{L}=-\sum_{\sigma}\psi_{\lambda}^{\dagger}\dot{\mathbf{k}}\cdot D^{k}_{i}\psi_{\lambda}.$ (33)
Expanding the wavefunction in momentum modes and replacing $\psi_{\lambda}(r)$ by $\psi_{\lambda}(k)$, one obtains—in the semiclassical approximation—the momentum-space Lagrangian for particle motion,
$\mathcal{L}_{semi}=\dot{\bm{k}}\cdot\bm{r}(k).$ (34)
Applying the Euler–Lagrange equations to this Lagrangian yields the semiclassical equation of motion
$\dot{\bm{r}}\sim\lambda\dot{\bm{k}}\times\bm{F},$ (35)
with $\bm{F}=\bm{\nabla_{k}}\times\bm{A}$. This relation provides the semiclassical description of the spin Hall effect of light [9, 8]. The dual gauge field, commonly identified with the Berry connection, has an associated Berry curvature that in momentum space takes the form
$\bm{F}=\lambda\mathbf{k}/k^{3},$ (36)
i.e., a “magnetic monopole” configuration [14, 5, 7]. These ideas have been applied to explain a variety of optical phenomena, including Rytov’s law  [30, 14], as well as spin-dependent transverse shifts such as the Imbert–Fedorov shift and the spin Hall effect of light  [19, 25, 35, 37, 33, 41, 10, 11, 17, 7, 12].
In the framework of noncommutative mechanics [26, 18, 16], the coordinate $r_{i}$ ceases to be commutative and instead obeys
$r_{i},r_{j}]=i\lambda\mathcal{F}_{ij}.$ (37)
This noncommutative structure is closely connected to the photon-coordinate algebra introduced by Schwinger in his analysis of the photon-localization problem [39, 38]. It therefore suggests that the issue of photon localization is fundamentally associated with the intrinsic dual gauge field.
By close analogy with an electron in an external electromagnetic field, the canonical momentum $\bm{P}$ and the mechanical momentum $\bm{p}$ are related through the vector potential as $\bm{P}-e\widetilde{\bm{A}}(\bm{r})=\bm{p}$. The associated momentum operators therefore obey the nonvanishing commutation relation
$p_{i},p_{j}]=ie\epsilon_{ijk}B^{k},$ (38)
with $\bm{B}=\bm{\nabla}\times\widetilde{\bm{A}}$. In the presence of a magnetic monopole of charge $g_{m}$, for which $\bm{B}=g_{m}\bm{r}/r^{3}$, a dual correspondence arises between the helicity $\lambda$ and the electric and magnetic charges [38].
$U_{g}=\exp[i\lambda\oint_{\Gamma}dk^{i}A_{i}(k)]=\exp[i\lambda\oint_{\Sigma}dS^{ij}\mathcal{F}_{ij}(k)],$ (39)
III Gauge field representation of spin-orbit coupling
In this section, we will examine how the spatial rotation of the dipole term $\bm{q}$ along the helical trajectories of photons induces nontrivial gauge fields, and show that these dipole-driven effects constitute the physical origin of the gauge fields associated with spin–orbit coupling.
We focus only on the Lagrangian of the dual gauge field in coordinate space. In the helicity-eigenstate representation $\psi_{\lambda}$, the effective Lagrangian associated with the gauge-field component $A_{s}$ is
@@ III Gauge field representation of spin-orbit coupling
Since photons can also carry intrinsic orbital angular momentum (IOAM), $\hat{\bf L}=\hat{\bf x}\times\hat{\bf k}$, we generalize the photon wave function to a twisted optical field, written as $\psi^{+}e^{il\varphi}$ and $\psi^{-}e^{-il\varphi}$, where $l$ is the topological charge and $\varphi$ is the azimuthal angle  [2].For such a twisted beam, an additional first-order correction arises from $\mathbf{q}\cdot\hat{\mathbf{k}}_{\perp}=\hat{L}_{z}\tau$ in $\Delta\hat{m}_{0}$, which yields
@@
which, for a vortex beam with topological charge $l$, becomes $\lambda_{\sigma}=(l+\sigma)$. This expression shows that, although the interactions of polarized light carrying topological charge $l$ with the medium appear formally on an equal footing [7], their underlying physical mechanisms are fundamentally different. In particular, the contribution associated with spin-orbit coupling originates from duality symmetry.
@@ IV Conclusion
We investigate the localization of the dual symmetry of the electromagnetic field in helical systems using a field-theoretical framework, in which a novel gauge field emerges naturally as a consequence of the local symmetry. We find that this dual gauge field renders the coordinates of the photon non-commutative, thereby revealing the fundamental nature of the gauge field responsible for the localization of photons: it serves as the bridge connecting the canonical coordinate and the physical coordinate. The spin–orbit interaction arising during beam propagation along helical trajectories is formally equivalent to the coupling between the helicity and the dual gauge field. This correspondence provides a unified gauge-theoretic interpretation of the spin Hall effect of light, where such phenomena can be viewed as analogues of axion–electromagnetic coupling [45, 23, 36, 44]. For vortex beams, corrections from orbit–orbit interactions remain significant, yet these are fundamentally distinct from spin–orbit coupling, which originates from a deeper intrinsic symmetry. This distinction provides a fresh perspective for differentiating between the spin and intrinsic orbital angular momentum (IOAM) of light.
A particularly important result is Eq. (40). In conventional electrodynamics, it is well known that charge conservation arises naturally from the $U(1)$ gauge symmetry of the electronic system, and that promoting this symmetry to a local one introduces the electromagnetic gauge potential—enabling interactions between electrons via photon exchange. In a completely analogous manner, the dual symmetry considered here is also a $U(1)$ symmetry, whose associated conserved quantity is helicity. Localizing this dual symmetry introduces a dual gauge field, implying that photons interact with each other through the exchange of quanta associated with this field. Consequently, our framework may provide a fundamentally new theoretical basis for describing interactions between light and matter.